Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 285     Entry time: 02 Aug 2023, 06:44
Author: Giuseppe Mercurio, Teguh Asmara 
Group: SCS 
Subject: shift summary 

achievements:

  • beam transport with the second half of undulators closed
  • measured pulse energy in SCS hutch and verified that the transmission of the beamline with hr grating is as previously reported
  • knife edge scans along vertical direction to find focus (<=11 um) and horizontal direction to have 175 um
  • half clip scans to find center of rotation
  • aligned Cu and Ni ML (theta offset) to find signal of Cu and Ni at hrixs
  • delivered the beam at the hrixs ccd camera

problems:

  • all the undulator cells were closed instead of only the second half of the undulators

plan:

  • optimize energy resolution at hrixs at Ni L edge (priority)

 

ELOG V3.1.4-7c3fd00